# Akash Srivastava2,3, Hongdong Li9, Junzhou Huang10, & Dimitris N. Metaxas1

### v2

@@ 3.2 Discrete Inversion for Controllable Editing
Max. The third way is inspired by the property of Gumbel distribution [54], that if $G_{1}$, $G_{2}$ are iid random variables following $\text{Gumbel}(\mu,\beta)$ then $\max{\{G_{1},G_{2}\}}-\beta\log{2}$ follows the same distribution. We also consider the max function for noise injection:
@@
We empirically find that linear strategy gives best results.
@@
Note that here we use the Gumbel max trick [27], which is equivalent to sampling from categorical distribution $q(x_{t}|x_{0})$.
@@
Note that here the latent $\bm{z}_{t}\in\mathbb{R}^{D\times K}$.
@@ 4 Experiments
In this section, we demonstrate the effectiveness of our proposed inversion methods on both image and language diffusion models. Our experiments show that the methods can preserve identity in both vision and language tasks while successfully making the intended changes. The implementation details can be reviewed in Supplementary Materials.
@@ 4.1 Image diffusion model
For the image diffusion model, we mainly investigate the use of absorbing state discrete model [2] including a masked generative model, Paella, and a multinomial diffusion model, VQ-Diffusion. We demonstrate the inversion reconstruction ability and image editing performance in both categories with DICE.
Dataset. The Prompt-based Image Editing Benchmark (PIE-Bench) by  [28] is a recently introduced dataset designed to evaluate text-to-image (T2I) editing methods. The dataset assesses language-guided image editing in 9 different scenarios with 700 images. The benchmark’s detailed annotations and variety of editing tasks were instrumental in thoroughly assessing our method’s capabilities, ensuring a fair and consistent comparison with existing approaches.
Figure 3: Visualization of editing results. Editing results for our method using Paella and VQ-Diffusion are presented, along with their corresponding prompts. The results demonstrate that our method can effectively modify the input image according to the target prompt while preserving the image structure. Editing with masked generative model (Paella [43]) is more stable and easier than with multinomial diffusion models (VQ-Diffusion [16]).
@@ 4.1.1 Inversion Reconstruction
Quantitative Analysis. The reconstruction performance of our method, as shown in Table 1, far surpasses the baseline Inpainting + Paella model across all metrics. In the case of masked inpainting, all image tokens are replaced with randomly sampled tokens, meaning the model lacks any prior information about the original image. As a result, the reconstructed image differs significantly from the one being inverted, leading to lower similarity scores. In contrast, our method demonstrates near-perfect reconstruction, as indicated by the metrics, and notably produces an identical image without the errors typically introduced by the VQ-VAE/GAN quantization process, as seen in the results marked with (). This highlights the superior accuracy and consistency of our approach in generating high-fidelity reconstructions.
@@
Discrete Inpainting + Paella Prompt 91.10 25.36 23.42
@@
Table 2: Quantitative results on image editing performance. Comparison of our proposed method with the masked inpainting with the discrete diffusion model Paella, as well as continuous diffusion model (Stable Diffusion v1.4) using DDIM inversion. “P2P” refers to Prompt-to-Prompt [21], and “Prompt” denotes editing performed solely through forward edit prompts. Entries marked with an asterisk () are cited from [28]. : For VQ-Diffusion, the images are down-sampled to $256\times 256$. It is important to note that due to differences in base models and editing algorithms, the metrics across methods are not directly comparable. However, our method significantly outperforms both inpainting and strong baselines (e.g., Null-Text Inversion + SD1.4) in terms of structural preservation. As expected, inpainting achieves a high CLIP score since it directly generates image patches based on the target prompt.
4.1.2 Editing Performance
Evaluation Metrics. To demonstrate the effectiveness and efficiency of our proposed inversion method, we employ eight metrics covering three key aspects: structure distance, background preservation, and edit prompt-image consistency, as outlined in [28]. We utilize the structure distance metric proposed by [51] to measure the structural similarity between the original and generated images. To evaluate how well the background is preserved outside the annotated editing mask, we use Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) [58], Mean Squared Error (MSE), and Structural Similarity Index Measure (SSIM) [53]. We also assess the consistency between the edit prompt and the generated image using CLIP [41] Similarity Score  [55], which is calculated over the whole image and specifically within the regions defined by the editing mask.
In Figure 3, we show the editing results for both Paella and VQ-Diffusion using DICE. Both models successfully modify real images according to the target prompts. In all cases, our results exhibit both high fidelity to the input image and adherence to the target prompt.
@@ 4.1.2 Editing Performance
Inversion+Model Editing PSNR $\uparrow$ LPIPS×${}_{{}^{\times 10^{3}}}$downarrow MSE×${}_{{}^{\times 10^{4}}}$downarrow SSIM×${}_{{}^{\times 10^{2}}}$uparrow
@@ 4.2 Language Diffusion Model
In this section, we evaluate DICE on RoBERTa [31], a text discrete diffusion model, to generate sentences with opposing sentiments while preserving structural similarities. We begin with two prompts—one with a positive sentiment and another with a negative sentiment. Each prompt contains two sentences: the first sentence indicates the sentiment type and sets the contextual background, and the second sentence is the target for inversion and generation. Initially, we invert the second sentence of the negative sentiment prompt using the entire prompt as context, which produces a noised token representation of that sentence. Next, we condition the model on the positive sentiment by concatenating the first sentence of the positive sentiment prompt with the noised token of the inverted negative sentence. This setup guides the model to generate a new second sentence that mirrors the structure of the original negative sentence but expresses a positive sentiment instead. Through this process, we assess the model’s capability to invert and generate text that aligns with a specified sentiment while retaining the original sentence’s structural elements.
@@
Negative Prompt Our Edited Results
It was dull and confusing. It was clear and surprising.
The project ended in failure. New project still in progress.
It’s complicated and not useful. It’s On and It’s Epic.
@@ 4.2.1 Inversion Reconstruction
Evaluation Metric. For reconstruction, we use Hit Rate, which is defined as the proportion of cases where each method generates an identical sentence to the original. In addition, we compute the Semantic Textual Similarity (STS) score by measuring the cosine similarity between the sentence embeddings, using the model proposed by [44].
@@
Editing Method Accuracy×${}_{{}^{\times 10^{2}}}$uparrow Textual Similarity×${}_{{}^{\times 10^{2}}}$uparrow
@@
Table 5: Text Inversion Reconstruction Performance. Quantitative comparisons of the text reconstruction performance by Masked Generation and DICE method using RoBERTa as the language model.
Editing Method Structure Preservation×${}_{{}^{\times 10^{2}}}$uparrow Sentiment Correctness×${}_{{}^{\times 10^{2}}}$uparrow
Masked Generation 29.80 12.94
Ours 94.76 72.51
@@ 4.2.2 Sentence Editing
In this section, we evaluate the editing performance of the proposed inversion method on RoBERTa.
Evaluation Metric. For the sentence editing task, we evaluate the generated sentences based on two criteria: (1) structural preservation, which assesses whether the sentence structure is retained, and (2) sentiment correctness, which evaluates whether the sentiment of the edited sentence aligns with the sentiment of the original prompt. Both the structural preservation rate and sentiment correctness rate are calculated using ChatGPT-4 [1] as a classifier. Qualitative samples are given in Table 7.
Results. Table 6 presents a comparative analysis of two text editing methods that both employ RoBERTa, focusing on the effectiveness in terms of Structure Preservation and Sentiment Correctness. Our method significantly outperforms masked generation in both metrics. This difference highlights the superior capability of our inversion method to encode the original structure of the text in the latent space and the flexibility to adjust its sentiment more accurately. In Supplementary Materials, we demonstrate both the initial prompt and the edited result. Our approach retains the sentence structure of the negative prompt while modifying its sentiment to a more positive one.
@@ 5 Conclusion
Acknowledgment. This research has been partially funded by research grants to Dimitris N. Metaxas through NSF: 2310966, 2235405, 2212301, 2003874, FA9550-23-1-0417, and NIH-5R01HL127661.
Appendix
Figure 4: This paper was accepted to CVPR 2025 but later desk-rejected post camera-ready.
@@ Appendix C Analysis on Mutual Information
Given a simple Gaussian DDPM with ${\bm{x}}_{0}\sim\mathcal{N}({\bm{0}},{\bm{I}})$, latents $\{{\bm{z}}_{t}\}$ are obtained with DDPM inversion [26], then the mutual information between ${\bm{z}}_{t}$ and ${\bm{x}}_{0}$:
@@
which is a Gaussian error term independent to $\bm{x}_{0}$ with mean 00 and variance $1-\overline{\alpha}_{t-1}+\alpha_{t}(1-\overline{\alpha}_{t})$. Thus we can calculate the mutual information
@@
We also provide the relationship between the mutual information of $z_{t},z_{0}$ and the timestep $t$ in Fig. 5.
@@ Appendix D Implementation Details
The hyper-parameters for Paella editing experiment is CFG$=10.0$, $\lambda_{1}=0.7$, $\lambda_{2}=0.3$ and $\tau=0.9$. The hyper-parameters for VQ-Diffusion in editing is CFG$=5.0$, $\lambda_{1}=0.2$, $\lambda_{2}=0.8$.
@@ E.1 Noise Injection Function
Max. The third way is inspired by the property of Gumbel distribution [54], that if $G_{1}$, $G_{2}$ are iid random variables following $\text{Gumbel}(\mu,\beta)$ then $\max{\{G_{1},G_{2}\}}-\beta\log{2}$ follows the same distribution. We also consider the max function for noise injection:
@@ Appendix F Additional Results on Image Editing
Additional baselines. We compare with SDEdit [35] and ControlNet [57] [^1] We use the ControlNet-InPaint model based on Stable Diffusion v1.5: https://github.com/mikonvergence/ControlNetInpaint. Results are shown in Figure 13 and Table 8.
@@
Appendix G Latent Space Properties
Figure 9: Latent space properties of DICE.
Figure 9 demonstrates additional experiment results on the latent space property including flipping and shifting in the latent space.
Appendix H Details on Text Editing Experiments
@@
Appendix I Additional Results on Sentiment Editing
@@
Table 8: Additional baselines. We compare with SDEdit [35] and ControlNet [57].

### v3

@@ 3.2 Discrete Inversion for Controllable Editing
Max. The third way is inspired by the property of Gumbel distribution (wikipedia_gumbel_distribution, ), that if $G_{1}$, $G_{2}$ are iid random variables following $\text{Gumbel}(\mu,\beta)$ then $\max{\{G_{1},G_{2}\}}-\beta\log{2}$ follows the same distribution. We also consider the max function for noise injection:
@@
We empirically find that linear strategy gives best results. The emperical studies can be find in Supplementary Materials Figure 8.
@@
Note that here we use the Gumbel max trick (jang2016categorical, ), which is equivalent to sampling from categorical distribution $q(x_{t}|x_{0})$. Note that below the latent $\bm{z}_{t}\in\mathbb{R}^{D\times K}$.
@@ 4 Experiments
In this section, we demonstrate the effectiveness of our proposed inversion methods on both image and language diffusion models. Our experiments show that the methods can preserve identity in both vision and language tasks while successfully making the intended changes. The implementation details are in Supplementary Materials Section D.
@@ 4.1 Image diffusion model
For the image diffusion model, we mainly investigate the use of absorbing state discrete model (austin2021structured, ) including a masked generative model, Paella, and a multinomial diffusion model, VQ-Diffusion. We show the inversion reconstruction and image editing performance in both categories with DICE.
Dataset. The Prompt-based Image Editing Benchmark (PIE-Bench) by  (ju2023direct, ) is a recently introduced dataset designed to evaluate text-to-image (T2I) editing methods. The dataset assesses language-guided image editing in 9 different scenarios with 700 images. The benchmark’s detailed annotations and variety of editing tasks were instrumental in thoroughly assessing our method’s capabilities, ensuring a fair and consistent comparison with existing approaches.
Table 1: Inversion Reconstruction performance. The metric is calculated between the original and inverted images. Due to the encoding and decoding steps in the VQ-VAE/GAN process, some inaccuracies are introduced by the quantization. $\dagger$ The PSNR is Inf due to the reconstruction of our method yielding the same VQ-VAE/GAN latents. Base model is Paella (rampas2022novel, ).
[…]
  Inpainting   10.50   565.11   1002.09   30.13
  Ours   30.91   39.81   11.07   90.22
  $\text{Ours}^{\dagger}$   Inf   0.07   0.01   99.99
Figure 3: Visualization of editing results. Editing results for our method using Paella and VQ-Diffusion are presented, along with their corresponding prompts. The results demonstrate that our method can effectively modify the input image according to the target prompt while preserving the image structure. Editing with masked generative model (Paella (rampas2022novel, )) is more stable and easier than with multinomial diffusion models (VQ-Diffusion (gu2022vector, )).
@@ 4.1.1 Inversion Reconstruction
Quantitative Analysis. The reconstruction performance of our method, as shown in Table 1, far surpasses the baseline Inpainting + Paella model across all metrics. In the case of masked inpainting, all image tokens are replaced with randomly sampled tokens, meaning the model lacks any prior information about the original image. As a result, the reconstructed image differs significantly from the one being inverted, leading to lower similarity scores. In contrast, our method demonstrates near-perfect reconstruction, as indicated by the metrics, and notably produces an identical image without the errors typically introduced by the VQ-VAE/GAN quantization process, as seen in the results marked with (). This highlights the superior accuracy and consistency of our approach in generating high-fidelity reconstructions. Visual results can be viewed in Figure 9 of the Supplementary Materials.
4.1.2 Editing Performance
Evaluation Metrics. To demonstrate the effectiveness of our proposed inversion method, we employ eight metrics covering three key aspects: structure distance, background preservation, and edit prompt-image consistency, as outlined in ju2023direct . We utilize the structure distance metric proposed by tumanyan2023plug to measure the structural similarity between the original and generated images. To evaluate how well the background is preserved outside the annotated editing mask, we use Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) (zhang2018unreasonable, ), Mean Squared Error (MSE), and Structural Similarity Index Measure (SSIM) (wang2004image, ). We also assess the consistency between the edit prompt and the generated image using CLIP (radford2021learning, ) Similarity Score  (wu2021godiva, ), which is calculated over the whole image and specifically within the regions defined by the editing mask.
Table 2: Quantitative results on image editing performance. Comparison of our proposed method with the masked inpainting with Paella, as well as continuous diffusion model (Stable Diffusion v1.4) using DDIM inversion. “P2P” refers to Prompt-to-Prompt (hertz2022prompt, ), and “Prompt” denotes editing performed solely through forward edit prompts. Entries marked with an asterisk () are cited from ju2023direct . : For VQ-Diffusion, the images are down-sampled to $256\times 256$. Please note that due to differences in base models and editing algorithms, the metrics across methods are not directly comparable. However, our method significantly outperforms both inpainting and strong baselines (e.g., Null-Text Inversion + SD1.4) in terms of structural preservation. As expected, inpainting achieves a high CLIP score since it directly generates image patches based on the target prompt.
@@ 4.1.2 Editing Performance
Discrete
Inpainting + Paella Prompt 91.10 25.36 23.42
@@
In Figure 3, we show the editing results for both Paella and VQ-Diffusion using DICE. Both models successfully modify real images according to the target prompts. In all cases, our results exhibit both high fidelity to the input image and adherence to the target prompt. Additional visualization results can be viewed in Figure 9 and 10 in Supplementary Materials.
@@
[…]
@@ 4.2 Language Diffusion Model
In this section, we evaluate DICE on RoBERTa (liu2019roberta, ) and LLaDA (nie2025large, ), a text discrete diffusion model, to generate sentences with opposing sentiments while preserving structural similarities. We begin with two prompts, one with a positive sentiment and another with a negative sentiment. Each prompt contains two sentences: the first sentence indicates the sentiment type and sets the contextual background, and the second sentence is the target for inversion and generation. Initially, we invert the second sentence of the negative sentiment prompt using the entire prompt as context, which produces a noised token representation of that sentence. Next, we condition the model on the positive sentiment by concatenating the first sentence of the positive sentiment prompt with the noised token of the inverted negative sentence. This setup guides the model to generate a new second sentence that mirrors the structure of the original negative sentence but expresses a positive sentiment instead. Through this process, we assess the model’s capability to invert and generate text that aligns with a specified sentiment while retaining the original sentence’s structural elements.
@@ 4.2.1 Inversion Reconstruction
Negative Prompt Our Edited Results
Negative: Regarding the lecture. Positive: Regarding the lecture.
It was dull and confusing. It was clear and surprising.
Negative: Despite the initial problems. Positive: Despite the initial problems.
The project ended in failure. New project still in progress.
Negative: Regarding the new app. Positive: Regarding the new app.
It’s complicated and not useful. It’s On and It’s Epic.
Evaluation Metric. For reconstruction, we use Hit Rate, which is defined as the proportion of cases where each method generates an identical sentence to the original. In addition, we compute the Semantic Textual Similarity (STS) score by measuring the cosine similarity between the sentence embeddings, using the model proposed by reimers2019sentence .
@@
Table 5: Text Inversion Reconstruction Performance. Evaluation of the text reconstruction performance by Masked Generation and DICE method using RoBERTa as the language model.
[…]
Textual
[…]
@@
Editing Method
Structure
[…]
Sentiment
[…]
Masked Generation + RoBERTa 29.80 12.94
Masked Generation + LLaDA 22.88 21.18
Ours + RoBERTa 94.76 72.51
Ours + LLaDA 94.12 72.29
@@ 4.2.2 Sentence Editing
In this section, we evaluate the editing performance of the proposed inversion method on RoBERTa and LLaDA.
Evaluation Metric. We evaluate the sentence sentiment editing task based on two criteria: (1) structural preservation, which assesses whether the sentence structure is retained, and (2) sentiment correctness, which evaluates whether the sentiment of the edited sentence aligns with the sentiment of the original prompt. Both the structural preservation rate and sentiment correctness rate are calculated using ChatGPT-4 (achiam2023gpt, ) as a classifier. Qualitative samples are given in Table 7.
Results. Table 6 presents a comparative analysis of two text editing methods that both employ RoBERTa and LLaDA, focusing on the effectiveness in terms of Structure Preservation and Sentiment Correctness. Our method significantly outperforms masked generation in both metrics. This difference highlights the superior capability of our inversion method to encode the original structure of the text in the latent space and the flexibility to adjust its sentiment more accurately. In Supplementary Materials, we demonstrate both the initial prompt and the edited result. Our approach retains the sentence structure of the negative prompt while modifying its sentiment to a more positive one.
@@ Appendix C Analysis on Mutual Information
Given a simple Gaussian DDPM with ${\bm{x}}_{0}\sim\mathcal{N}({\bm{0}},{\bm{I}})$, latents $\{{\bm{z}}_{t}\}$ are obtained with DDPM inversion (huberman2024edit, ), then the mutual information between ${\bm{z}}_{t}$ and ${\bm{x}}_{0}$:
@@
which is a Gaussian error term independent to $\bm{x}_{0}$ with mean 0 and variance $1-\overline{\alpha}_{t-1}+\alpha_{t}(1-\overline{\alpha}_{t})$. Thus we can calculate the mutual information
@@
We also provide the relationship between the mutual information of $z_{t},z_{0}$ and the timestep $t$ in Figure 5.
@@ Appendix D Implementation Details
The hyper-parameters for Paella editing experiment is CFG=10.0=10.0, $\lambda_{1}=0.7$, $\lambda_{2}=0.3$ and $\tau=0.9$. The hyper-parameters for VQ-Diffusion in editing is CFG=5.0=5.0, $\lambda_{1}=0.2$, $\lambda_{2}=0.8$.
@@ E.1 Noise Injection Function
Max. The third way is inspired by the property of Gumbel distribution (wikipedia_gumbel_distribution, ), that if $G_{1}$, $G_{2}$ are iid random variables following $\text{Gumbel}(\mu,\beta)$ then $\max{\{G_{1},G_{2}\}}-\beta\log{2}$ follows the same distribution. We also consider the max function for noise injection:
@@ Appendix F Additional Results on Image Editing
Additional baselines. We compare with SDEdit meng2021sdedit and ControlNet zhang2023adding  [^1] We use the ControlNet-InPaint model based on Stable Diffusion v1.5: https://github.com/mikonvergence/ControlNetInpaint. Results are shown in Figure 12 and Table 8.
@@
Appendix G Details on Text Editing Experiments
@@
Appendix H Additional Results on Sentiment Editing
@@
Table 8: Additional baselines. We compare with SDEdit meng2021sdedit and ControlNet zhang2023adding .
@@
Appendix I Limitations
While Discrete Inversion shows promise, we empirically find that editing with multinomial diffusion models may not work as robustly as with masked generative models. Furthermore, it may appear less effective in style transfer tasks, such as transforming an image of a cat into a silver cat statue. Interesting future directions include: (1) developing a more theoretical analysis of mutual information and convergence for continuous and discrete inversion algorithms, (2) extending Discrete Inversion to score distillation sampling, and (3) exploring the integration of Semantic Guidance within discrete settings.